OpenStudy (ashley644):

How many mL of 2.0M hydrochloric acid is required to neutralize 100. ML of 0.80 M potassium hydroxide?

OpenStudy (photon336):

well @ashley644 we can use this formula: \[M_{a}V_{a} = M_{b}V_{b}\] the little a means acid: so it's molarity of the acid times the volume of the acid = the molarity of the base times the volume of the base. taking a deeper look at this formula, something interesting happens here. what this formula is saying is that the number of moles of acid is equal to the number of moles of base when we have a neutralization reaction. \[M_{a}V_{a} = Molarity*volume = \frac{ moles }{ L }*(L) = moles \] First ask yourself what you are looking for:

OpenStudy (photon336):

Be sure to convert mL to L by multiplying by dividing by 1,000 100mL/1000 = 0.1 L of 0.8 M potassium hydroxide this is our base 2.0 M HCL our acid. \[\frac{ M_{base}*V_{base} }{ M_{acid} } = Volume~of~acid \] \[\frac{ 0.8M*0.1L }{ 2.0M } = Volume_{HCL}\]